### v1

# Bi-CoG: Bi-Consistency-Guided Self-Training for Vision-Language Models

Abstract
Exploiting unlabeled data through semi-supervised learning (SSL) or leveraging pre-trained models via fine-tuning are two prevailing paradigms for addressing label-scarce scenarios. Recently, growing attention has been given to combining fine-tuning of pre-trained vision-language models (VLMs) with SSL, forming the emerging paradigm of semi-supervised fine-tuning. However, existing methods often suffer from model bias and hyperparameter sensitivity, due to reliance on prediction consistency or pre-defined confidence thresholds. To address these limitations, we propose a simple yet effective plug-and-play methodology named Bi-Consistency-Guided Self-Training (Bi-CoG), which assigns high-quality and low-bias pseudo-labels, by simultaneously exploiting inter-model and intra-model consistency, along with an error-aware dynamic pseudo-label assignment strategy. Both theoretical analysis and extensive experiments over 14 datasets demonstrate the effectiveness of Bi-CoG, which consistently and significantly improves the performance of existing methods.
Introduction
Addressing downstream tasks with limited labeled data remains a critical challenge in real-world applications (Yang et al. 2022). Two prominent paradigms have emerged to tackle this issue: (1) adapting pre-trained models, such as vision-language models (VLMs) (Radford et al. 2021; Jia et al. 2021), for image classification tasks; and (2) applying semi-supervised learning (SSL) methods that leverage large amounts of inexpensive unlabeled data. Recent work (Lv et al. 2025b) has further suggested combining these two approaches, where the strong zero-shot capabilities of pre-trained VLMs are used to generate reliable pseudo-labels that guide semi-supervised fine-tuning, thereby fully exploiting the strengths of both paradigms.
Figure 1: Comparison of existing pseudo-labeling paradigms and our proposed Bi-CoG. (a) Majority voting from multiple VLMs; (b) Top-$k$ selection based on predicted logits; (c) Bi-CoG dynamically generate accurate and unbiased pseudo-labels based on bi-consistency and model error.
Recent efforts have explored integrating these two paradigms by leveraging pre-trained models to extract domain knowledge from unlabeled data for downstream task adaptation. A common approach is model ensembling (Figure 1.a), such as V-PET (Gupte et al. 2024), which aggregates multi-view one-hot predictions from multiple vision models. As shown in Figure 1.b, another line of work selects top-$k$ predictions using class-wise or confidence-based thresholds. For instance, GRIP (Menghini, Delworth, and Bach 2023) and UPL (Huang, Chu, and Wei 2022) enforce class balance by selecting equal samples per class; CPL (Zhang et al. 2024) refines labels via instance-level confidence matrices to reduce pre-training bias.
However, these methods struggle to balance the trade-off between pseudo-label accuracy and bias. Ensemble-based approaches can produce highly accurate pseudo-labels, but they often inadvertently amplify biases inherited from pre-training, leading to uneven performance across subgroups—a phenomenon commonly referred to as the “Matthew effect” (Zhu, Luo, and Liu 2022; Chen et al. 2022; Gan and Wei 2024). In contrast, soft-labeling techniques and class-wise pseudo-label allocation strategies can alleviate overconfidence to some extent, but they tend to introduce additional label noise and are highly sensitive to manually tuned hyperparameters. To the best of our knowledge, adaptively generating reliable and unbiased pseudo-labels in response to model performance remains a critical yet underexplored challenge in semi-supervised fine-tuning.
To address this challenge, we propose a novel method named Bi-Consistency-Guided Self-Training (Bi-CoG). As illustrated in Figure 1.c, Bi-CoG leverages multiple VLMs as base learners to perform dynamic self-training guided by both inter- and intra-model consistency as well as model error estimation. Specifically, we conduct a dynamic self-training process that comprises three-stage pseudo-label selection. In the first stage, we apply majority voting to generate ensembled pseudo-labels, ensuring reliable supervision. Then, we assess each model’s bias on a sample using weak and strong augmentations: predictions consistent under no and weak augmentation but changing under strong augmentation indicate reliability and less prone to pre-training bias. Finally, based on theoretical insights, we dynamically adjust the upper bound on the number of pseudo-labels used for training, according to models’ relative accuracy on labeled data. This further filters pseudo-labels and enables adaptive self-training throughout the training process.
As a plug-and-play approach, Bi-CoG can be seamlessly integrated into existing VLM fine-tuning pipelines. To evaluate its effectiveness, we conduct experiments under both open-world generalization (Zhou et al. 2024) and standard semi-supervised learning(Yang et al. 2022) settings, combining Bi-CoG with various prompt-tuning methods. Extensive results demonstrate the effectiveness of our bi-consistency-guided dynamic self-training strategy, which consistently improves performance across 14 datasets without relying on larger models or external knowledge, outperforming state-of-the-art (SOTA) methods.
Our main contributions are summarized as follows:
  • We propose Bi-CoG, a plug-and-play method that jointly leverages the pre-trained knowledge of VLMs and external information from unlabeled data.
  • We design a deeply coupled self-training strategy integrates inter-model consistency, intra-model consistency and model error estimation, enabling reliable, low-bias pseudo-label generation and adaptive training.
  • Theoretical analysis and extensive experiments demonstrate that Bi-CoG consistently improves performance, achieving up to a 5.69% gain across 14 datasets and surpassing SOTA methods.
Related Work
Parameter-efficient Fine-tuning on VLMs
While pre-trained VLMs demonstrate strong zero-shot generalization in domain-specific downstream tasks, their large parameter sizes often hinder efficient adaptation to specific applications (Zhou et al. 2022b). To address this, recent works have proposed parameter-efficient fine-tuning approaches that introduce a small number of trainable parameters while keeping the majority of the model frozen (Li et al. 2024). Tip-Adapter (Zhang et al. 2022) and CLIP-Adapter (Gao et al. 2024) append lightweight bottleneck layers after the text or image encoders to learn residual features for better representation. CoOp (Zhou et al. 2022b) introduces prompt tuning to VLMs for the first time by replacing manually crafted hard prompts (e.g., “a photo of a classname”) with learnable soft text prompts optimized via backpropagation. Building on this idea, methods such as MaPLe (Khattak et al. 2023a), PromptSRC (Khattak et al. 2023b), and BMIP (Lv et al. 2025a) extend prompt tuning to both visual and textual encoders, enabling joint vision-language adaptation. While these methods have advanced the application of VLMs to downstream tasks, the transformation of prompts into continuous vectors makes them prone to overfitting and leads to sub-optimal performance in open-world scenarios (Zhou et al. 2022a; Khattak et al. 2023b).
Pseudo-labeling in Semi-supervised Learning
Semi-supervised learning (SSL) addresses label scarcity by leveraging both labeled and unlabeled data. Pseudo-labeling, the most widely used SSL approach, assigns labels to unlabeled data based on model predictions and improves performance through self-training. The key to effective pseudo-labeling lies in the generation of high-quality pseudo-labels (Yang et al. 2022). Traditional SSL methods typically filter pseudo-labels using confidence thresholds or adopt data augmentation to generate diverse views, aiming to improve label quality and model robustness (Lee et al. 2013a). For VLMs, their strong generalization ability allows them to produce highly accurate pseudo-labels even in zero- or few-shot settings, making them strong candidates as pseudo-labelers. Recent works have explored more advanced strategies: GRIP (Menghini, Delworth, and Bach 2023) gradually increases the upper limit on the number of pseudo-labels to mitigate calibration errors (Oh et al. 2024) and label imbalance (Wang et al. 2022); and CPL (Zhang et al. 2024) progressively refines pseudo-labels based on a confidence score matrix to avoid overconfidence. Despite their effectiveness, these methods struggle with balancing pseudo-label noise and inherited pre-training biases (Oh et al. 2024; Wang et al. 2022), while failing to dynamically adapt to model performance, ultimately limiting their potential.
Figure 2: The overall framework of Bi-CoG’s pseudo-label selection. During the self-training phase, Bi-CoG employs inter-model consistency, intra-model consistency, and error-aware filtering to generate accurate and unbiased pseudo-labels, which are then used to fine-tune the target VLM.
Preliminary
In this section, we first revisit the core principles of CLIP (Radford et al. 2021), one of the most widely used VLMs, along with its associated prompt tuning methods, and subsequently formalize our problem definition.
CLIP.
CLIP comprises a text encoder $\psi$ and an image encoder $\phi$, aligned via pretraining on massive image-text pairs. Given an image $x$ and a set of class labels $\mathcal{Y}={y_{1},\dots,y_{C}}$, CLIP classifies $x$ by computing the similarity between the image representation $z=\phi(x)$ and each class representation $w_{i}=\psi(\pi_{i})$, where $\pi_{i}$ is a natural language prompt such as "a photo of a [CLASSi]". The predicted probability is computed as follows:
$p(y=i\mid x)=\frac{\exp(\cos(\phi(x),\psi(w_{i}))/\tau)}{\sum_{j=1}^{C}\exp(\cos(\phi(x),\psi(w_{j}))/\tau)}$ (1)
where $\tau$ is a temperature parameter that controls the sharpness of the probability distribution.
Prompt Tuning.
As a representative PEFT method for CLIP-based VLMs, prompt tuning learns adaptable prompts instead of fine-tuning the text encoder $\psi$. For instance, some studies (Zhou et al. 2022b, a) replace manually designed prompt templates with learnable prompt vectors $P=[t_{1},t_{2},\dots,t_{n}]$, where each $t_{i}$ is a trainable token vector and $n$ denotes the prompt length. Given a set of class names $c_{1},c_{2},\dots,c_{N}]$, the prediction probability is computed as:
$p(y=i\mid x)=\frac{\exp(\cos(\phi(x),\psi([P,c_{i}]))/\tau)}{\sum_{j=1}^{N}\exp(\cos(\phi(x),\psi([P,c_{j}]))/\tau)}$ (2)
where $\cdot$$\cdot$] denotes concatenation. The prompt $P$ is optimized via backpropagation during fine-tuning.
Problem Definition.
In downstream scenarios with scarce labeled data, we typically have access to a small labeled dataset denoted as $D_{L}=\{(x_{i},y_{i})\}_{i=1}^{M}$, along with a large amount of unlabeled data $D_{UL}=\{x_{i}\}_{i=1}^{N}$, where $M\ll N$. Given a collection of $K$ pre-trained VLMs $\{f_{k}\}_{k=1}^{K}$ (with $K\geq 3$), the objective is to effectively leverage both $D_{L}$ and $D_{UL}$ to fine-tune the models and maximize performance on the target dataset $D_{test}$.
Methodology
To address the prevalent challenges of balancing pseudo-label accuracy with model bias and reducing the strong reliance on empirical parameter tuning in prior methods, we propose Bi-Consistency-Guided Self-Training (Bi-CoG), a simple yet effective plug-and-play framework. As illustrated in Figure 2, the Bi-CoG framework comprises the following three major components for pseudo-label selection:
  • A model ensembling module that aggregates predictions through majority voting based on inter-model consistency, generating high-quality candidate pseudo-labels to ensure reliable supervision.
  • A multi-view filtering mechanism that leverages intra-model consistency by applying diverse transformations to unlabeled data and discarding biased samples with inconsistent predictions, thereby mitigating the influence of pre-training bias.
  • An error-aware bounding strategy that dynamically estimates model performance and sets an upper limit on the number of pseudo-labels used, promoting stable and progressive self-training.
Algorithm 1 Bi-CoG’s overall learning algorithm.
Input: Pre-trained VLMs $\{f_{k}\}_{k=1}^{K}$, labeled data $D_{L}$, unlabeled data $D_{UL}$, scaling factor $\alpha$
Output: Fine-tuned VLMs
[…]
2:  $e^{\prime}_{j}\leftarrow 0.5$, $\mathrm{update}_{j}\leftarrow\mathrm{False}$
3:end for
[…]
5:while $\mathrm{improve}$ do
6:  $t\leftarrow t+1$
[…]
8:   $\mathrm{update}_{j}\leftarrow\mathrm{False}$
9:   $e_{j}\leftarrow\mathrm{MeasureError}(D_{L},\{f_{k}\}_{\ k\neq j}^{K})$
10:   if $e_{j}<e_{j}^{\prime}$ then
11:    $D_{Inter}^{j}\leftarrow\mathrm{InterConsistency}(D_{UL},\{f_{k}\}_{k\neq j}^{K})$
12:    $D_{Intra}^{j}\leftarrow\mathrm{IntraConsistency}(D_{UL},\{f_{k}\}_{k\neq j}^{K})$
13:    $\mathrm{update}_{j}\leftarrow\mathrm{ErrorAware}\left(e_{j},\ e_{j}^{\prime},\ t\right)$
14:    $D_{PL}^{j}\leftarrow\mathrm{Subsample}(D_{Inter}^{j}\cap D_{Intra}^{j},(\frac{e_{j}^{\prime}}{e_{j}})^{\alpha t}L_{j}^{\prime})$
15:   end if
16:   if $\mathrm{update}_{j}$ then
17:    $e_{j}^{\prime}\leftarrow e_{j}$, $f_{j}\leftarrow\mathrm{FineTune}(f_{j},D_{L}\cup D_{PL}^{j})$
18:   end if
19:  end for
20:  if $\mathrm{update}==\left[\mathrm{False}\right]*K$ then
21:   $\mathrm{improve}\leftarrow\mathrm{False}$
22:  end if
23:end while
24:return $\{f_{k}\}_{k=1}^{K}$
The main procedure of Bi-CoG is summarized in Algorithm 1, and we elaborate on each component in the subsequent sections.
Inter-model Consistency
During self-training with pseudo-labeled data, label fidelity is critical to effective and stable optimization (Angluin and Laird 1988). Excessive noise can mislead learning, causing convergence to suboptimal or incorrect representations. To mitigate this issue, we first generate pseudo-labels based on inter-model consistency by integrating predictions of multiple VLMs, with the primary goal of ensuring the accuracy of pseudo-labels.
Concretely, for each model $f_{j}$ ($1\leq j\leq K$), we construct a candidate pseudo-labeled dataset $D_{Inter}^{j}$ by selecting unlabeled instances on which the majority of the other K1K-1 models reach a consensus. For each unlabeled sample $x_{i}\in D_{UL}$, we first compute the class-wise vote count from the other models as follows:
$v_{i}(y)=\sum_{1\leq k\leq K,\ k\neq j}\mathbb{I}\left[f_{k}(x_{i})=y\right]$ (3)
If the most voted label $y^{*}=\arg\max_{y\in\mathcal{Y}}v_{i}(y)$ satisfies $v_{i}(y^{*})\geq\left\lceil\frac{K-1}{2}\right\rceil$, we assign $x_{i}$ a pseudo-label $\hat{y}_{i}=y^{*}$ and include the pair $(x_{i},y^{*})$ in $D_{Inter}^{j}$. Otherwise, the sample is discarded. Formally, this candidate set is defined as:
$D_{Inter}^{j}=\left\{(x_{i},y^{*})\ \middle|\ v_{i}(y^{*})\geq\left\lceil\frac{K-1}{2}\right\rceil\right\}$ (4)
where $\mathcal{Y}$ denotes the label space, and $\mathbb{I}[\cdot]$ is the indicator function. This inter-model consistency strategy ensembles multiple models to generate more reliable pseudo-labels, reducing noise from individual prediction errors. Moreover, the leave-one-out scheme prevent amplifying biases inherent to the target model, thereby resulting in more accurate and unbiased candidate datasets for fine-tuning.
Intra-model Consistency
While inter-model consistency improves pseudo-label reliability by aggregating agreement across models, it may still preserve instances where VLMs share similar biases, often due to overlapping pre-training corpora or architectural homogeneity (Gan and Wei 2024; Wang and Russakovsky 2023). To further mitigate biased supervision, we introduce intra-model consistency, which leverages the consistency of a single model’s predictions under different input views to refine the pseudo-labeled sets obtained in the previous stage.
Specifically, for each unlabeled sample $x_{i}\in D_{UL}$, we apply both weak and strong data augmentations. Each model $f_{k}$ then produces predictions on the original input $x_{i}$, the weakly augmented $\tilde{x}_{i}^{w}$, and the strongly augmented $\tilde{x}_{i}^{s}$. A sample is retained only if the prediction remains consistent between the original and weakly augmented inputs, i.e., $f_{k}(x_{i})=f_{k}(\tilde{x}_{i}^{w})$, but differs under strong augmentation, i.e., $f_{k}(\tilde{x}_{i}^{s})\neq f_{k}(x_{i})$. This criterion identifies samples with low uncertainty and meaningful sensitivity, as evidenced by prediction consistency under weak perturbations and inconsistency under strong perturbations. The intra-model consistent pseudo-label set for $f_{j}$ is defined as the intersection of predictions from the other models:
$D_{Intra}^{j}=\bigcap_{\begin{subarray}{c}k=1\\ k\neq j\end{subarray}}^{K}\left\{(x_{i},f_{k}(x_{i}))\middle|\begin{array}[]{l}f_{k}(x_{i})=f_{k}(\tilde{x}_{i}^{w})\\ f_{k}(x_{i})\neq f_{k}(\tilde{x}_{i}^{s})\end{array}\right\}$ (5)
Once the $D_{Intra}^{j}$ are computed, the candidate pseudo-label set for target model $f_{j}$ is constructed by intersecting its inter-model consistent set $D_{Inter}^{j}$ with the intra-model consistent set :
$D_{PL}^{j}=D_{Inter}^{j}\cap D_{Intra}^{j}$ (6)
The intersection of pseudo-labels from inter- and intra-model consistency preserves accuracy while mitigating shared model biases. Consequently, the resulting pseudo-labeled datasets are both highly reliable and low in bias, providing a robust foundation for the subsequent dynamic self-training stage.
Error-aware Filtering
After bi-consistency-guided pseudo-labeling, we obtain a candidate sample set $D_{PL}$, which contrains high-accuracy and low-biased pseudo-labels. However, during self-training, it is crucial to determine an appropriate upper bound on the number of pseudo-labeled samples based on model performance. Prior studies (Angluin and Laird 1988; Chen et al. 2022) have shown that even minor label noise, if not aligned with model competence, can significantly degrade performance, making further noise reduction in $D_{PL}$ essential. Existing methods (Menghini, Delworth, and Bach 2023; Huang, Chu, and Wei 2022) typically rely on fixed thresholds to limit the number of pseudo-labels, which may be suboptimal as model performance evolves over time.
To address this limitation, we dynamically adjust the upper bound on pseudo-label usage based on estimated model error ratio, guided by a theoretical analysis. Specifically, following prior work (Angluin and Laird 1988; Zhou and Li 2005) we formalize the influence of noisy supervision on model $f_{j}$ during iterative self-training:
Lemma 1.
Let $L_{t}$ denote the training set $D_{train}$ for model $f_{j}$ at iteration $t$, and $e_{t}$ represent the relative error rate of the remaining K1K-1 models at the same iteration, i.e., the probability that the majority vote produces an incorrect label. Then, training $f_{j}$ on $D_{train}$ will result in performance improvement if the following condition is satisfied:
$0<\frac{e_{t}}{e_{t-1}}<\frac{L_{t-1}}{L_{t}}<1$ (7)
According to Lemma 1, the model will be improved throughout the training process as long as Equation 7 holds.
Average ImageNet Caltech101 OxfordPets StanfordCars
Method HM Acc. HM Acc. HM Acc. HM Acc. HM Acc.
CLIP 73.00 66.41 70.18 66.70 95.68 93.30 94.11 89.10 68.85 65.60
CoOp 72.10 66.59 72.00±0.34 68.73±0.25 90.91±1.61 89.90±0.99 93.97±0.72 88.73±0.09 69.94±0.59 66.40±0.50
+ Bi-CoG 77.79 71.03 73.06±0.08 69.78±0.04 96.67±0.05 94.24±0.25 96.28±0.60 89.57±2.02 73.38±0.31 69.31±0.76
MaPLe 78.78 72.18 73.58±0.09 70.30±0.14 96.58±0.21 94.87±0.09 96.60±0.36 92.33±0.31 73.53±0.81 69.97±0.87
+ Bi-CoG 79.77 73.54 74.11±0.07 70.82±0.09 96.77±0.17 95.09±0.46 96.93±0.06 92.56±0.12 74.20±0.31 70.48±0.36
PromptSRC 79.99 74.11 74.07±0.06 70.77±0.05 96.11±0.08 94.63±0.12 96.26±0.29 92.13±0.59 76.68±0.16 73.03±0.19
+ Bi-CoG 81.46 74.96 74.18±0.03 70.94±0.03 96.45±0.16 94.74±0.16 96.37±0.13 92.62±0.13 79.07±0.15 73.77±0.01
Flowers102 Food101 FGVCAircraft SUN397 DTD
Method HM Acc. HM Acc. HM Acc. HM Acc. HM Acc.
CLIP 74.44 70.70 90.39 85.60 31.21 24.80 72.37 62.60 56.62 44.00
CoOp 75.70±2.82 71.27±2.21 85.82±1.41 80.03±1.17 29.93±1.54 26.07±0.74 71.60±1.81 63.33±1.44 55.93±1.25 50.87±0.62
+ Bi-CoG 83.57±0.91 79.59±1.22 90.54±0.14 85.03±0.65 35.61±0.29 28.40±0.21 76.89±0.06 67.15±0.18 59.90±1.37 50.96±1.35
MaPLe 82.98±0.19 78.47±0.60 90.83±0.35 86.47±0.40 35.30±0.88 27.30±0.96 79.42±0.30 70.97±0.24 67.40±3.09 55.30±1.40
+ Bi-CoG 84.62±0.25 79.12±0.51 91.67±0.03 87.46±0.02 35.58±0.71 27.63±0.56 80.27±0.19 71.82±0.23 69.88±1.07 56.36±0.73
PromptSRC 86.13±0.39 80.77±0.50 90.91±0.12 86.47±0.12 34.99±6.68 29.93±2.50 80.68±0.23 72.13±0.26 70.72±1.23 58.87±1.28
+ Bi-CoG 87.60±0.18 83.19±0.55 91.25±0.01 86.65±0.02 39.89±0.34 31.98±0.69 80.89±0.04 72.43±0.10 71.67±0.42 59.48±0.10
EuroSAT UCF101 CIFAR-10 CIFAR-100 ImageNet-100
Method HM Acc. HM Acc. HM Acc. HM Acc. HM Acc.
CLIP 60.21 48.30 74.42 67.50 88.43 79.00 55.17 46.00 88.33 86.50
CoOp 72.04±3.29 59.50±2.10 67.03±1.48 64.23±1.26 89.17±1.32 77.00±2.42 53.42±3.23 45.03±2.88 81.97±2.38 81.17±1.96
+ Bi-CoG 85.84±1.98 67.98±0.67 73.11±1.99 68.01±0.92 92.07±0.46 82.28±0.99 61.71±0.26 52.55±0.32 90.39±0.35 89.59±0.30
MaPLe 78.35±1.46 60.90±8.71 81.57±0.39 74.70±0.57 92.02±0.76 83.83±1.21 64.73±0.30 56.10±0.37 90.00±0.17 89.07±0.33
+ Bi-CoG 81.33±2.89 70.52±2.81 82.23±0.58 75.75±0.51 92.15±0.09 84.41±0.69 66.68±0.52 57.97±0.61 90.41±0.29 89.52±0.32
PromptSRC 81.87±2.52 70.03±0.86 82.30±0.87 76.67±0.37 91.46±0.30 83.80±0.42 67.23±0.42 58.93±0.39 90.49±0.22 89.43±0.26
+ Bi-CoG 86.28±0.87 71.01±0.47 83.09±0.24 77.30±0.44 94.50±0.15 84.96±1.75 68.19±0.38 60.20±0.07 91.07±0.02 90.15±0.05
Table 1: Performance comparison on 14 datasets using ViT-B/16 architecture. The best performance is in bold.
In practice, we estimate the theoretical error rate ratio from the performance of the K1K-1 models on the labeled set. However, this estimation can be biased, as repeated training on labeled data tends to produce stable—and often low—error ratios, which are reflected in Equation 7 as values close to 1. In contrast, the actual error ratio on unlabeled data typically decreases during training, making the true ratio $\frac{\hat{e}_{t}}{\hat{e}_{t-1}}$ lower than the estimated $\frac{e_{t}}{e_{t-1}}$. Ignoring this discrepancy would introduce bias into the estimated pseudo-label budget, compromising the validity of the theoretical analysis. To address this issue, we introduce Theorem 1 to provide a correction:
Theorem 1.
As the training progresses, the true error rate ratio can be approximately equal to the $t$-power of the estimated error rate ratio, as shown below:
$\frac{e_{t}}{e_{t-1}}\approx\left[\frac{\hat{e}_{t}}{\hat{e}_{t-1}}\right]^{\alpha t}$ (8)
where $\alpha>0$ is a scaling factor that reflects the error reduction rate. Then, the original condition in Equation 7 is approximately equivalent to the following constraints:
$L_{t}$ $\leq\left\lceil\left(\frac{\hat{e}_{t-1}}{\hat{e}_{t}}\right)^{\alpha t}\times L_{t-1}-1\right\rceil$ (9)
$L_{t-1}$ $>\frac{\hat{e}_{t}^{\alpha t}}{\hat{e}_{t-1}^{\alpha t}-\hat{e}_{t}^{\alpha t}}$
A detailed proof of Lemma 1 and Theorem 1 is provided in the supplementary material. Based on the theoretical analysis, we establish the necessary condition under which the model can be effectively optimized by self-training, as described in Equation 9, and establish an upper bound on the pseudo-label budget for each iteration. Accordingly, the final set of pseudo-labeled samples is formalized as:
$D_{PL^{*}}^{j}=\mathrm{Subsample}\left(D_{PL}^{j},\ \left(\frac{e_{j}^{\prime}}{e_{j}}\right)^{\alpha t}L_{j}^{\prime}\right)$ (10)
This dynamic filtering strategy reduces the mismatch between model capacity and pseudo-label quantity. Consequently, the model can self-regulate the amount of pseudo-labeled data during training, leading to stable and consistent performance improvements.
In summary, Bi-CoG’ offers the following advantages:
  1. 1.
    Plug-and-Play Adaptability: Bi-CoG can be seamlessly integrated into existing pretrain-finetune pipelines, providing consistent performance improvements with minimal changes to the original framework.
  2. 2.
    Robust Pseudo-Labeling: By jointly leveraging bi-consistency and error-aware mechanisms, Bi-CoG generates more reliable pseudo-labels, effectively reducing label noise and alleviating model bias.
  3. 3.
    Adaptive Self-Training Control: Empowered by the error-aware mechanism, Bi-CoG can automatically regulate the self-training process, eliminating the need for manually preset iteration counts and enhances adaptability across different scenarios.
Experiment
In this section, we design experiments to answer the following research questions:
  1. 1.
    RQ1: Does Bi-CoG enhance existing PEFT methods for VLMs in open-world generalization scenarios?
  2. 2.
    RQ2: Can Bi-CoG deliver competitive performance compared to conventional SSL approaches?
  3. 3.
    RQ3: How does Bi-CoG compare to other methods that jointly leverage pre-trained models and unlabeled data?
  4. 4.
    RQ4: What is the contribution of each component within the Bi-CoG framework?
Experimental Settings
Evaluation Paradigm.
To answer RQ1 and RQ2, we evaluate the effectiveness of Bi-CoG on two widely adopted settings: open-world generation setting (Zhou et al. 2024; Lv et al. 2025a) and semi-supervised setting (Lee et al. 2013b; Yang et al. 2022). In open-world generation setting, the model is trained only on a subset of categories (referred to as base classes), while the remaining categories are treated as novel classes. We report the harmonic mean (HM) of the model’s performance on base and novel classes, as well as the overall accuracy on the full test set, which includes both base and novel classes. Semi-supervised setting follows a common protocol where only a small portion of the data is labeled, while the majority remains unlabeled.
Dataset and Baselines.
We evaluate the performance of our method on 14 recognition datasets, including 11 commonly used for assessing VLMs and 3 classical benchmarks for semi-supervised learning. Detailed information about the datasets is provided in the supplementary material. We select CLIP (Radford et al. 2021) and its three representative prompt tuning methods as baselines: CoOp (Zhou et al. 2022b), MaPLe (Khattak et al. 2023a), and PromptSRC (Khattak et al. 2023b), among which PromptSRC represents the current SOTA method. For methods that integrate the two paradigms, we select GRIP (Menghini, Delworth, and Bach 2023) and CPL (Zhang et al. 2024) as baselines, with CPL representing the current SOTA approach. The parameters of Bi-CoG and additional experimental details are provided in the supplementary material.
Experimental Results
RQ1:
Does Bi-CoG enhance existing PEFT methods for VLMs in open-world generalization scenarios?
To evaluate the robustness of Bi-CoG in open-world scenarios, we conduct a comprehensive study by integrating it with three representative prompt tuning methods across 14 datasets. For each dataset, we report two metrics: HM of base and novel class accuracies, and overall accuracy (Acc.). We set the base-to-novel class ratio to 50:50 and report the mean and standard deviation over three runs. As shown in Table 1, Bi-CoG consistently improves both metrics across all datasets, achieving the highest average performance. Notably, when applied to CoOp, Bi-CoG yields over a 5% improvement in HM. Moreover, Bi-CoG demonstrates particularly strong gains on traditional open-world benchmarks. For example, on low-resolution datasets such as CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), it achieves up to an 8% increase in HM, underscoring its effectiveness across datasets with varying resolutions. These results highlight the effectiveness of Bi-CoG and emphasize the importance of integrating unsupervised strategies with pre-trained models for improved generalization in open-world settings.
RQ2:
Can Bi-CoG deliver competitive performance compared to conventional SSL approaches?
Method CIFAR-10 CIFAR-100
4 shots 25 shots 4 shots 25 shots
ReMixMatch 80.90±9.64 94.56±0.05 55.72±2.06 72.57±0.31
FixMatch (RA) 86.19±3.37 94.93±0.65 51.15±1.75 71.71±0.11
FixMatch (CTA) 88.61±3.35 94.93±0.33 50.05±3.01 71.36±0.24
CoOp 80.60±1.79 85.10±0.29 49.00±0.83 57.43±0.25
+ Bi-CoG 84.90±1.06 86.10±0.03 52.77±0.24 58.05±0.05
PromptSRC 86.03±0.48 90.10±0.08 60.33±0.17 66.33±0.29
+ Bi-CoG 89.83±0.35 91.07±0.42 60.56±0.79 66.84±0.67
Table 2: Performance on low-resolution SSL datasets.
As shown in Table 2, SSL algorithms exhibit a significant advantage over baseline prompt tuning methods on SSL datasets. This is largely due to the lower resolution of these datasets, which makes it more difficult for VLMs to effectively capture visual content (Krizhevsky, Hinton et al. 2009; Lv et al. 2025b). This limitation is also reflected in the sensitivity to the number of shots: increasing the number of shots from 4 to 25 yields much larger performance gains for SSL algorithms compared to VLM-based prompt tuning. However, Bi-CoG provides the model with more reliable training examples, enabling prompt tuning methods to achieve SOTA performance under the challenging 4-shot setting.
These results demonstrate that Bi-CoG can effectively leverage unlabeled data to capture richer visual representations, even in low-resolution and data-scarce scenarios.
RQ3:
How does Bi-CoG compare to other methods that jointly leverage pre-trained models and unlabeled data?
To further validate the effectiveness of Bi-CoG, we conducted comparative experiments against the current SOTA methods GRIP (Menghini, Delworth, and Bach 2023) and CPL (Zhang et al. 2024) on the Flowers102 dataset (Nilsback and Zisserman 2008). Following GRIP, we adopt the same base-to-novel class ratio of 62:38, and use CoOp as the underlying prompt tuning method for Bi-CoG and other approaches to ensure faireness. As shown in Table 3, our Bi-CoG achieves the best performance under both settings, outperforming CPL by up to 5.9 percentage points in SSL setting. More importantly, Bi-CoG eliminates the need for manually setting hyperparameters such as the number of pseudo-labels $k$, effectively decreasing the reliance on manual tuning and susceptibility to internal model bias.
In summary, Bi-CoG represents one of the most effective integrations of pretrain-finetuning and SSL methods to date.
Ablation Study
To answer RQ4, we first examine the impact of the bi-consistency-guided pseudo-labeling strategy on both pseudo-label accuracy and the distribution of biased samples. As shown in Figure 3.a, removing the inter-model consistency module results in a notable decline in pseudo-label accuracy, highlighting the importance of ensemble-based labeling for generating high-quality supervision. Notably, although the bi-consistency strategy exhibits relatively lower pseudo-label accuracy in the early stages of training, its accuracy steadily improves as training progresses. In contrast, other single-strategy pseudo-labeling methods show either a decline or fluctuation in accuracy, demonstrating the stability and reliability of bi-consistency-guided pseudo-labeling. Furthermore, as illustrated in Figure 3.b, incorporating intra-model consistency helps smooth the class-wise distribution of pseudo-labels, effectively mitigating the bias inherited from pre-training. These results collectively confirm that the bi-consistency-guided strategy produces pseudo-labels that are both more accurate and less biased—two critical factors for stable and effective self-training.
Figure 3: (a) Pseudo-label accuracy over training iterations on the DTD dataset under different pseudo-labeling strategies; (b) Distribution of pseudo-labels on the EuroSAT dataset before and after applying the intra-model consistency.
Method Domain Data SSL Open-World
CLIP Zero-shot 72.08 77.80
FPL Few-shot + Unlabeled 75.96 80.97
IFPL 78.68 82.08
GRIP 83.60 86.26
CPL 89.66 87.35
Bi-CoG 96.45 88.01
$\Delta$ +6.79 +0.66
Table 3: Performance comparison on the Flowers102 dataset, with a base-to-novel class ratio of 62:38. Our method outperforms prior approaches under both settings.
Figure 4: Comparison between the estimated and true error rate ratios on StanfordCars under the open-world setting. The scaling factor is fixed at $\alpha=1$ throughout training.
Method Open-World
HM Acc.
CoOp + Bi-CoG 77.79 71.03
- w/o. Inter-model consistency 76.14 68.83
- w/o. Intra-model consistency 75.85 69.08
- w/o. Error-aware filtering 76.92 70.92
Table 4: Ablation study of Bi-CoG’s pseudo-labeling strategy under the open-world generalization setting. Experimental results show that each component plays a critical role in improving model performance.
To further validate the theoretical foundation of our error-aware mechanism, we analyze the evolution of the actual error rate ratio on pseudo-labeled samples and its estimation from labeled data on the StanfordCars dataset (Krause et al. 2013). As illustrated in Figure 4, the condition in Equation 8 holds reliably during training, supporting the practical validity of our theoretical derivation and confirming the soundness of the Bi-CoG’ dynamic self-training framework.
Finally, we conduct a comprehensive ablation study on the key components of Bi-CoG, using CoOp as the underlying prompt tuning method and reporting the average performance across 14 datasets. As shown in Table 4, any component of our three-stage pseudo-labeling strategy, which integrats bi-consistency and error-aware filtering, plays a critical role in improving model performance. This highlights the deeply coupled and effective design of the pseudo-labeling mechanism in Bi-CoG.
Conclusion
This paper investigates the integration of pre-trained vision-language models with semi-supervised learning for downstream tasks with sparse labeled data. We identify several critical issues in existing approaches, including the trade-off between label quality and model bias, as well as the limited adaptability arising from the reliance on manually set fixed thresholds. To address these challenges, we propose Bi-CoG, a plug-and-play self-training framework that incorporates a three-stage, deeply coupled pseudo-labeling strategy with adaptive adjustment capabilities. Theoretical analysis and experimental results across 14 datasets in both open-world generalization and semi-supervised settings demonstrate that Bi-CoG can be seamlessly integrated with any pretrain-finetuning methods, consistently achieving performance improvements.
Supplementary Material
A. Detailed Theoretical Analysis
A.1. Proof of Lemma 1
Proof.
When the aggregated training data contains noisy samples, let $\mathcal{F}$ denote the hypothesis space, $\eta$ the noise ratio in the training set ($\eta<0.5$), $\epsilon>0$ the target error rate, and $\delta<1$ the confidence parameter. According to PAC learning theory (Angluin and Laird 1988; Zhou and Li 2005), if the number of training samples $m$ satisfies
$m\geq\frac{2}{\epsilon^{2}(1-2\eta)^{2}}\ln\left(\frac{2|\mathcal{F}|}{\delta}\right),$ (11)
then, for the learned model $f$ and the optimal model $f^{*}$ within $\mathcal{F}$, we have
$\mathbb{P}\left[d(f,f^{*})\geq\epsilon\right]\leq\delta,$ (12)
where $d(f,f^{*})$ denotes the excess error rate of $f$ over $f^{*}$.
Let $c=2\mu\ln\left(\frac{2|\mathcal{F}|}{\delta}\right)$, where $\mu$ makes is Eq. (11) hold equality. Eq. (11) can then be rewritten as
$m=\frac{c}{\epsilon^{2}(1-2\eta)^{2}}.$ (13)
Thus, the relationship between the error rate $\epsilon$, the number of training samples $m$, and the noise ratio $\eta$ is given by
$\epsilon=\frac{\sqrt{c}}{\sqrt{m(1-2\eta)^{2}}}.$ (14)
In the Bi-CoG training framework, we maintain $K$ vision-language models (VLMs). Suppose our optimization target is the $j$-th model, denoted as $f_{j}$. We adopt the leave-one-out scheme and generate a pseudo-label candidate set $D_{PL}$ using the remaining K1K-1 models via inter-model consistency and intra-model consistency. At iteration $t$, the candidate training set is
$D_{train}=D_{L}\cup D_{PL},$ (15)
Let $e_{t}$ be the probability that the prediction obtained by a majority vote (agreed upon by more than half of the K1K-1 models) is still incorrect. Denote $|D_{L}|=L$ and $|D_{PL}|=L_{t}$, and approximate $|D_{\text{train}}|\approx L+L_{t}$. The noise ratio at iteration $t$ is then estimated as
$\eta_{t}=\frac{e_{t}\times L_{t}}{L+L_{t}}.$ (16)
From Eq. (14), $\epsilon$ is inversely proportional to $\sqrt{m(1-2\eta)^{2}}$. Therefore, the condition $\epsilon_{t}<\epsilon_{t-1}$ is equivalent to
$(L+L_{t})$ $\left(1-2\times\frac{e_{t}\times L_{t}}{L+L_{t}}\right)^{2}>$ (17)
$(L+L_{t-1})$ $\left(1-2\times\frac{e_{t-1}\times L_{t-1}}{L+L_{t-1}}\right)^{2}$
A sufficient condition for $\epsilon_{t}<\epsilon_{t-1}$ is
$\begin{cases}L_{t}>L_{t-1},\\ e_{t}\times L_{t}<e_{t-1}\times L_{t-1},\end{cases}$ (18)
which can be equivalently expressed as
$0<\frac{e_{t}}{e_{t-1}}<\frac{L_{t-1}}{L_{t}}<1.$ (19)
Hence, we complete the proof. ∎
A.2. Proof of Theorem 1
Proof.
Let $e_{t}$ denote the true relative error rate of pseudo-labels at iteration $t$, and let $\hat{e}_{t}$ denote its estimated value computed on the labeled dataset at the same iteration. We begin by assuming that, during the course of training, the ratio of the true error rates between two consecutive iterations can be closely approximated by a power-law function of the corresponding ratio of the estimated error rates, given by:
$\frac{e_{t}}{e_{t-1}}\approx\left(\frac{\hat{e}_{t}}{\hat{e}_{t-1}}\right)^{\alpha t}$ (20)
where $\alpha$ is a scaling factor characterizing the rate of error reduction. Substituting this relationship into the condition in Eq. (19), we obtain
$0<\left(\frac{\hat{e}_{t}}{\hat{e}_{t-1}}\right)^{\alpha t}<\frac{L_{t-1}}{L_{t}}<1$ (21)
Dataset Domain Categories Train Test
ImageNet Natural picture 1,000 1.28M 50,000
Caltech101 Object picture 100 4,128 2,465
OxfordPets Pet picture 37 2,944 3,669
StanfordCars Car picture 196 6,509 8,041
Flowers102 Flower picture 102 4,093 2,463
Food101 Food picture 101 50,500 30,300
FGVCAircraft Aircraft picture 100 3,334 3,333
SUN397 Scene picture 397 15,880 19,850
DTD Texture picture 47 2,820 1,692
EuroSAT Satellite picture 10 13,500 8,100
UCF101 Video frame 101 7,639 3,783
CIFAR-10 Natural picture 10 50,000 10,000
CIFAR-100 Natural picture 100 50,000 10,000
ImagetNet-100 Natural picture 100 130,000 5,000
Table 5: Details on the datasets used in our experiments.
Therefore, the maximum allowable number of pseudo-labeled samples at iteration $t$ is given by
$L_{t}^{*}=\left\lceil\left(\frac{\hat{e}_{t-1}}{\hat{e}_{t}}\right)^{\alpha t}\times L_{t-1}-1\right\rceil$ (22)
In addition, the number of pseudo-labeled samples at iteration t1t-1 must satisfy
$L_{t-1}<\left\lceil\left(\frac{\hat{e}_{t-1}}{\hat{e}_{t}}\right)^{\alpha t}\times L_{t-1}-1\right\rceil$ (23)
which is equivalent to
$L_{t-1}>\frac{\hat{e}_{t}^{\alpha t}}{\hat{e}_{t-1}^{\alpha t}-\hat{e}_{t}^{\alpha t}}$ (24)
Consequently, we obtain both the upper bound on the number of pseudo-labeled samples at iteration $t$ and the corresponding lower bound constraint on $L_{t-1}$. ∎
B. Implementation Details
Transformation Description Parameter Value
Weak Augmentations
RandomHorizontalFlip Flips the image horizontally with a probability of $p$. pp $0.5$
RandomCrop Crops a random patch of the image after padding. size (224, 224)
padding 28
Strong Augmentations
ColorJitter Randomly changes the brightness, contrast, saturation, and hue. brightness 0.4
contrast 0.4
saturation 0.1
hue 0.4
RandomGrayscale Converts the image to grayscale with a probability of $p$. pp $0.2$
GaussianBlur Applies Gaussian blur to the image with a probability of $p$. kernel_size 21
pp 0.5
Table 6: Details of Weak and Strong Data Augmentations.
Dataset Details.
We evaluate our method on 14 recognition datasets, comprising 11 commonly used benchmarks for evaluating vision-language models (VLMs) and 3 classical benchmarks for semi-supervised learning. These datasets include CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), ImageNet and ImageNet-100 (Deng et al. 2009), Caltech101 (Fei-Fei, Fergus, and Perona 2004), OxfordPets (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), Food101 (Bossard, Guillaumin, and Van Gool 2014), Flowers102 (Nilsback and Zisserman 2008), FGVCAircraft (Maji et al. 2023), SUN397 (Xiao et al. 2010), UCF101 (Soomro, Zamir, and Shah 2012), DTD (Cimpoi et al. 2014), and EuroSAT (Helber et al. 2019).
For ImageNet, Caltech101, OxfordPets, StanfordCars, Flowers102, FGVCAircraft, SUN397, UCF101, DTD, and EuroSAT, we follow the dataset splits used in CoOp (Zhou et al. 2022a). For CIFAR-10, CIFAR-100, and ImageNet-100, we adopt the standard semi-supervised splits used in FixMatch (Sohn et al. 2020). Detailed statistics of these datasets are provided in Table 5. Specifically, for the open-world generalization experiment, we follow the standard base-to-novel setting(Zhou et al. 2022a), where base and novel classes are split evenly. We sample 16 shots per class as labeled data, and treat all remaining training samples as unlabeled. Due to computational constraints, for ImageNet we randomly sample 5% of the unlabeled set for training.
Training Details.
As a plug-and-play method, Bi-CoG can be seamlessly integrated into any pretrain–finetune paradigm. In our experiments, we adopt CoOp (Zhou et al. 2022b), MaPLe (Khattak et al. 2023a), and PromptSRC (Khattak et al. 2023b) as the base methods, keeping all hyperparameters identical to those reported in their original papers. Unless otherwise specified, we set the number of models to $K=3$ and the scaling factor to $\alpha=1$. In the dynamic self-training stage, each model is first warmed up for $\zeta$ epochs using a small set of labeled data. Then, we apply our three-stage pseudo-label selection strategy to obtain pseudo-labels with high accuracy and low model bias, which are further used for self-training over $\kappa$ additional epochs. The detailed parameters are provided in Table 7. During the intra-model consistency stage, we apply weak and strong augmentations to each sample, with the specific augmentation strategies summarized in Table 6. We conduct all experiments on a single NVIDIA A800 GPU.
Dataset CoOp MaPLe PromptSRC
$\zeta$ $\kappa$ $\zeta$ $\kappa$ $\zeta$ $\kappa$
CIFAR-100 15 1 5 1 15 1
ImageNet 15 3 5 1 15 1
FGVCAircraft 15 3 5 1 20 1
SUN397 15 3 5 1 15 1
UCF101 15 3 5 1 15 1
CIFAR-10 15 3 5 1 15 3
DTD 15 3 5 1 15 1
EuroSAT 15 3 7 1 5 20
StanfordCars 15 3 7 1 5 20
ImageNet-100 15 5 5 1 15 3
Caltech101 20 5 5 1 20 5
OxfordPets 20 5 7 1 15 1
Flowers102 20 5 5 1 20 5
Food101 20 5 5 1 15 1
Table 7: Details of the training epoch.